OpenStudy (anonymous):

Find the slope of the line using the two plotted points. (-7,3) and (5, -9)

OpenStudy (anonymous):

rise over run; -(3+9)/(7+5) =-12/12 =-1

OpenStudy (anonymous):

Whoops, forgot about those pesky negative numbers.

OpenStudy (anonymous):

Thanks Blakey